OpenStudy (anonymous):

Find the integral 2(e^(-.1t))dt

OpenStudy (lalaly):

\[\int\limits{2e^{-t}dt}\]\[=\frac{2e^{-t}}{-1}+C\]

OpenStudy (lalaly):

or is that 0.1t?

OpenStudy (anonymous1):

\[\int 2e^{-0.1t}dt\] Substitute u = -0.1t and du = -0.1 dt.

OpenStudy (lalaly):

yeah^^

OpenStudy (anonymous1):

Actually I'm not sure whether the asker meant -1 or -0.1.

OpenStudy (anonymous):

-0.1

OpenStudy (anonymous1):

In this case, substitute u = -0.1t and du = -0.1 dt. \[\frac{1}{-0.1}\int 2e^{u}du = -20\int e^{u}du\]

OpenStudy (anonymous1):

You should get: \[-20e^{-0.1t}+\text{constant}\]

OpenStudy (anonymous):

yes that is what I got thank you

OpenStudy (anonymous1):

You're welcome.